# RSCC: A Large-Scale Remote Sensing Change Caption Dataset for Disaster Events

### v4

@@ Abstract
Remote sensing is critical for disaster monitoring, yet existing datasets lack temporal image pairs and detailed textual annotations. While single-snapshot imagery dominates current resources, it fails to capture dynamic disaster impacts over time. To address this gap, we introduce the Remote Sensing Change Caption (RSCC) dataset, a large-scale benchmark comprising 62,315 pre-/post-disaster image pairs (spanning earthquakes, floods, wildfires, and more) paired with rich, human-like change captions. By bridging the temporal and semantic divide in remote sensing data, RSCC enables robust training and evaluation of vision-language models for disaster-aware bi-temporal understanding. Our results highlight RSCC’s ability to facilitate detailed disaster-related analysis, paving the way for more accurate, interpretable, and scalable vision-language applications in remote sensing. Code and dataset are available at https://github.com/Bili-Sakura/RSCC.
Figure 1: Examples of RSCC.
@@ 2.5 Post-Correction and Human Verification
To ensure the reliability of captions generated by QvQ-Max, we implement a two-stage post-correction process. First, the Qwen2.5-Max [54] systematically enforces metadata alignment by correcting discrepancies in disaster type (e.g., resolving mismatches between "hurricane" and metadata-specified "flooding") and damage descriptions (e.g., revising "minor damage" to "destroyed" based on building annotations). This automated stage achieves disaster type consistency with metadata from 98.7% to 100.0%. Second, a subset of RSCC captions (988 samples) is manually validated by three experts using a 0/1 binary rubric across four criteria: disaster type accuracy, damage detail completeness, factual consistency, and clarity. 92.3% of sampled captions passed validation, with 89.3% inter-annotator agreement (Cohen’s ϰ). Failed captions were reprocessed through the automated pipeline with refined rules, ensuring final dataset consistency. Full details of correction rules and evaluation protocols are provided in Appendix A.
@@ 3.2 RSCC for vision-language model training
In order to facilitate the vision-language model training, we divide the RSCC dataset into two splits, where a train split contains 61,327 image pairs from 31 distinct events across xBD as well as EBD and a test split contains 988 image pairs from 19 distinct events in xBD. We conduct full-parameter fine-tuning on Qwen2.5-VL 7B [57] using train set for 2 epoch with a batch_size=1batch\_size=1 on a single node equipped with 2 NVIDIA H800 GPUs. We initialize the learning rate at 1e-6 and 1e-5 for LLM backbone and vision encoder respectively and employ a cosine learning rate decay schedule for optimization. For image inputs, we maintain the native resolution of RSCC as 512×512 for maximum pixel inputs and minimum pixel inputs. The training procedure cost a total of 40 GPU clock hours.
@@ 5.2 Remote Sensing Change Caption Dataset
6 Limitations
Due to the lack of proficient labels and complexity of image pairs themselves, the generated captions may contains vague descriptions which is even hard for experts to clarify. Besides, we only employ text similarity metrics because existing image–to-text captioning metrics (e.g., FLEUR [28], SPARC [26] and G-VEval [67]) only focus on single image which fail in multi-image scenario. We leave these parts for future work.

### v5

@@ Abstract
Remote sensing is critical for disaster monitoring, yet existing datasets lack temporal image pairs and detailed textual annotations. While single-snapshot imagery dominates current resources, it fails to capture dynamic disaster impacts over time. To address this gap, we introduce the Remote Sensing Change Caption (RSCC) dataset, a large-scale benchmark comprising 62,351 pre-/post-disaster image pairs (spanning earthquakes, floods, wildfires, and more) paired with rich, human-like change captions. By bridging the temporal and semantic divide in remote sensing data, RSCC enables robust training and evaluation of vision-language models for disaster-aware bi-temporal understanding. Our results highlight RSCC’s ability to facilitate detailed disaster-related analysis, paving the way for more accurate, interpretable, and scalable vision-language applications in remote sensing. Code and dataset are available at https://github.com/Bili-Sakura/RSCC.
Figure 1: A sample from RSCC dataset.
@@ 2.5 Post-Correction and Human Verification
To ensure the reliability of captions generated by QvQ-Max, we implement a two-stage post-correction process. First, the Qwen2.5-Max [57] systematically enforces metadata alignment by correcting discrepancies in disaster type (e.g., resolving mismatches between "hurricane" and metadata-specified "flooding") and damage descriptions (e.g., revising "minor damage" to "destroyed" based on building annotations). This automated stage achieves disaster type consistency with metadata from 93.2% to 100.0%. Second, a subset of RSCC captions (10%) is randomly selected and manually validated by three experts using a 0/1 binary rubric across four criteria: disaster type accuracy, damage detail completeness, factual consistency, and clarity. 100.0% of sampled captions passed validation. Failed captions were reprocessed through the automated pipeline with refined rules, ensuring final dataset consistency. Full details of correction rules and evaluation protocols are provided in Appendix A.
@@ 3.2 RSCC for vision-language model training
In order to facilitate the vision-language model training, we divide the RSCC dataset into two splits, where a train split contains 61,363 image pairs from 31 distinct events across xBD as well as EBD and a test split contains 988 image pairs from 19 distinct events in xBD. We conduct full-parameter fine-tuning on Qwen2.5-VL 7B [58] using train set for 2 epoch with a batch_size=1batch\_size=1 on a single node equipped with 2 NVIDIA H800 GPUs. We initialize the learning rate at 1e-6 and 1e-5 for LLM backbone and vision encoder respectively and employ a cosine learning rate decay schedule for optimization. For image inputs, we maintain the native resolution of RSCC as $\times512$ for maximum pixel inputs and minimum pixel inputs. The training procedure cost a total of 40 GPU clockwall hours.
@@ 5.2 Remote Sensing Change Caption Dataset
6 Limitations and Discussions
Due to the lack of proficient labels and complexity of image pairs themselves, the generated captions may contains vague descriptions which is even hard for experts to clarify. Besides, we only employ text similarity metrics because existing image-to-text captioning metrics (e.g., FLEUR [30], SPARC [28] and G-VEval [68]) only focus on single image which fail in multi-image scenario. We leave these parts for future work.
Also, our preliminary study have tested baseline VLMs on change detection and multi-label classification upon RSCC, where it show that naively employing VLMs such tasks yields much inferior results compared to specialized models [83, 65, 14]. Moreover, a recent paper [17] denotes that VLMs’ performance would degrade compared to their visual encoders only. Thus, we generally recommend using specialized models for such visual-centric tasks, and we hope the community will develop strong VLMs that are able to naively deal with these tasks.
@@ 7 Conclusions
8 Acknowledgements
We gratefully acknowledge the valuable suggestions and preliminary work contributed by student members Tesi Lin and Zeyu Zhang. We also thank Alibaba Cloud (Aliyun) for providing access to the Qwen API, which was instrumental in this project.
This work was supported by the National Natural Science Foundation of China under Grants 42394060 and 42394062 as well as the National Key Research and Development Program of China under Grant 2019YFE0127400.
@@ A.2 More Results
Figure 13: RSCC (EBD + xBD) distribution. Image Credit: Wang et al. [77].